# Resolving UnderEdit & OverEdit with Iterative & Neighbor-Assisted Model Editing

### v1

@@ 5.2 Neighbor-Assisted Model Editing Results
To evaluate neighbor-assisted model editing method, we only conducted experiments on COUNTERFACT due to data limitations of ZsRE explained in Appendix B. We perform the iterative model editing with the modified neighbor loss function. We excluded the collapsed experimental settings. The evaluation results are shown in Table 2.
We observed consistently higher specificity across all iterations when using neighbor-assisted editing denoted by NA_ compared to setups without it. This shows the effectiveness of enforcing neighboring knowledge unchanged during the Optimization stage.
With increase in specificity, we observed an increase in overall score as well. Specifically, we observed an increase in success accuracy of up to 6 percentage points, with a greater improvement in specificity accuracy of up to 34 percentage points (NA_MEMIT on GPT-2 XL). Although there is a decrease in generalization, the improvement in efficacy is more significant, resulting in an overall increase in the score.
Moreover, the proposed stopping criteria $|\Delta p_{k}|\leq\epsilon$ defined for iterative model editing remains effective for neighbor-assisted model editing. We did observed an exception when Llama-2 (7B) neighbor-assisted edited with PMET showed a performance decrease instead of increase across all metrics. We attribute this degradation to increase in tendency to collapse indicated by high model perplexity measured by ME-PPL-50. Notably, Llama-2 (7B) was again the only model to exhibit this collapse behavior, reinforcing our earlier hypothesis that model-specific factors contribute to instability. However, pinpointing the training-related factors responsible for this behavior requires deeper analysis, which we leave for future work.
In conclusion, while both our proposed methods—iterative and neighbor-assisted model editing—do not inherently prevent collapse, they remain susceptible when other contributing factors are present.
6 Analysis
This section dives deeper into the proposed methods. Specifically, we aim to understand how iterative model editing addresses UnderEdit, the impact of the number of neighbors added in neighbor-assisted model editing for resolving OverEdit, and the effectiveness of the stopping criteria.
6.1 How does iterative model editing address UnderEdit?
We hypothesize that the iterative model editing approach can reduce the number of UnderEdit cases. To test this hypothesis, we identified UnderEdit examples in GPT-2 XL edited with MEMIT on COUNTERFACT after the first iteration, i.e, edits $(s_{i},r_{i},o_{i},o^{*}_{i})$ where $f_{\theta_{2}}(x_{i})[o_{i}^{*}]<f_{\theta_{2}}(x_{i})[o_{i}]$. We then tracked $|\Delta p_{k}|$ and efficacy accuracy across subsequent iterations. Figure 4 illustrates the results of the analysis. We observe that the rate of $|\Delta p_{k}|$ decrease and accuracy improving over iterations is much more pronounced for the UnderEdit examples. This observation tells us that multiple iterations of Spread is the larger contributor to getting higher performance.
Accuracy Accuracy Accuracy Accuracy
Unedited 0 1.00% 1.00%
MEMIT 4 99.00% 38.00% 52.00% 54.00%
NA_MEMIT 4 99.00% 36.00% 86.00% 60.00%
NAP_MEMIT 4 99.00% 37.00% 84.00% 61.00%
PMET 8 99.00% 31.00% 67.00% 52.00%
NA_PMET 9 98.00% 29.00% 85.00% 53.00%
NAP_PMET 8 98.00% 30.00% 84.00% 54.00%
6.2 How prefixes influence neighbor-assisted model editing behavior?
Using random prefixes aid generalization across contexts in the memory editing process when only the target knowledge edit is known. So we pose the question, does adding random prefixes to the neighbor knowledge prompts help prevent OverEdit? To answer this question we run an experiment by adding random prefixes to the neighboring knowledge used during edit. Table 3 shows an increase in specificity accuracy [^4] The full results with success on each measurement and perplexity performance are in Table 9. However, this increase is less compared to the improvement of using neighbor-assisted editing versus no neighbor-assist. Regardless, the prefix-free neighbor-assisted edits (NA_) achieved better overall performance denoted by Accuracy in Score due to their significant boost in specificity.
@@ 6.2 How prefixes influence neighbor-assisted model editing behavior?
6.3 How effective is the stopping criterion?
@@ 7 Related Work
In this work, we extensively discussed locate and edit model editing methods—ROME Meng et al. (2023a), MEMIT Meng et al. (2023b), and PMET Li et al. (2024). In addition, there is a body of research that employs meta-learners to guide the parameter updates required for specific edits. For example, KE Cao et al. (2021) uses a hyper-network to update model parameters, while MEND Mitchell et al. (2022a) trains gradient-based, lightweight model editor networks. MALMEN Tan et al. (2024) builds upon MEND to address scalability challenges.
Another line of research adds new knowledge without altering the model’s parameters. SERAC Mitchell et al. (2022b), GRACE Hartvigsen et al. (2023), and WISE Wang et al. (2024a) achieve this by employing additional memory to store new knowledge. A router network is then trained to decide whether to retrieve knowledge from the original model or the additional memory, ensuring the intended knowledge is accessed without modifying the model’s core parameters. We specifically focus on locate-and-edit model editing methods due to their effectiveness and efficiency in updating only the important parameters. Our proposed method introduces simple changes to existing techniques while still demonstrating effectiveness.
@@ 9 Limitations
Limited computational resources restricted us from experimenting with larger batch sizes and additional LLMs, such as GPT-NeoX (20B) and larger Llama-2 models. We hypothesize that the experimental result trend will remain the same, and we leave the verification of this hypothesis for future work.
Appendix A MEMIT vs PMET vs ROME
As discussed in Section 2, model editing algorithms—MEMIT Meng et al. (2023b), PMET Li et al. (2024), and ROME Meng et al. (2023a)—operate on the hypothesis that updating the final MLP parameters is sufficient to increase the likelihood of a new object $o^{*}$ over the original object $o$ when presented with a (subject, relation) pair $x=(s,r)$ as input to the LLM. Specifically, the final MLP weight matrix $W$ functions as a linear associative memory that stores a key-value mapping $[k,v]$ [^5] Distinct from key-value pairs in attention mechanismsMeng et al. (2023a). Here, the key encodes the last subject token, while the value represents the relation-object pair $(r,o)$ as a property of the subject $s$. Within the transformer architecture, the key corresponds to the output of the first fully connected MLP layer, whereas the value corresponds to the output of the second fully connected MLP layer $z$, as shown in Figure2. This key-value mapping $[k,v]$ is derived by computing the inner product between the key $k$ and the final MLP weight matrix $W$ as $Wk\approx v$.
@@ Appendix A MEMIT vs PMET vs ROME
MEMIT’s
objective in the Spread stage is to update final MLP weight matrix $W_{l}$ at each causal layer $l$ with a small change $\Delta$, producing new weights $W^{*}_{l}$ such that the $M$ new key-value mappings $[K_{M},V_{M}]$ are incorporated. Simultaneously, MEMIT seeks to preserve $E$ existing key-value mappings $[K_{E},V_{E}]$, where $K_{E}=[k_{e}]_{e=1}^{E}$ and $V_{E}=[v_{e}]_{e=1}^{E}$ are pre-existing vector keys and their corresponding values.
MEMIT obtains $W^{*}_{l}$ by solving an optimization problem that seek a transformation $\hat{W}_{l}$ minimizing the sum of squared euclidean distances between the transformed key vectors $\hat{W}_{l}k_{i}$ and their respective target vectors $v_{i}$:
$W_{1}$ $\triangleq\arg\min_{\hat{W}}\Bigg{(}\sum_{i=1}^{E}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}$
$\quad+\sum_{i=E+1}^{E+M}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}\Bigg{)}.$
The term $\sum_{i=E+1}^{E+M}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}$ enforces the modification of $M\gg 1$ knowledge entries, while $\sum_{i=1}^{E}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}$ iensures the preservation of existing knowledge. For a complete solution, we refer the reader to Meng et al. (2023b).
@@
follows the same optimization objective as MEMIT in the Spread stage but differs slightly in the Optimization stage. As illustrated in Figure 2, PMET searches for an ideal self-attention output $\bar{attn}^{l_{c}}_{t}$ and an ideal MLP output $\bar{z}^{l_{c}}_{t}$. It hypothesizes that the self-attention module encodes general knowledge patterns, and thus, its contribution to the hidden state $h^{l_{c}}_{t}$ is not required when editing specific knowledge in the MLP. Consequently, PMET’s Spread stage updates the final MLP weights $W^{l}$ in the causal layers using only the ideal MLP output $\bar{z}^{l_{c}}_{t}$. This modification allows PMET to achieve more precise model edits compared to MEMIT, leading to improved efficacy and generalization performance.
@@
is the predecessor of MEMIT. Unlike MEMIT and PMET, ROME assumes that any single causal layer can sufficiently store the new knowledge. Thus, in the Optimization stage, it identifies the ideal MLP output $\bar{z}^{l_{*}}_{t}$ at a single arbitrary causal layer $l_{*}$ and subsequently updates only the final MLP weights $W^{l_{*}}$ at that layer in the Spread stage, as shown in Figure 2.
ROME partially follows MEMIT’s optimization objective but with a stricter constraint: it aims to compute the updated MLP weight $\hat{W}^{l_{*}}$ such that it preserves all existing $E$ key-value mappings $[K_{E},V_{E}]$ while inserting only a single new key-value mapping ($k_{*},v_{*}$). Meng et al. (2023a) achieves this via the following closed-form solution:
minimize $\text{minimize }||\hat{W}^{l_{*}}K_{E}-V_{E}||$
such that $\text{such that }\hat{W}^{l_{*}}k_{*}=v_{*}$
For further details and the complete derivation, we refer the reader to Meng et al. (2023a).
ROME, due to its strict equality constraint, does not support batch editing [^6] Since ROME can only edit one fact at a time, batch edits are performed sequentially.. Notably, ROME produces the most precise edits compared to MEMIT and PMET when modifying a single fact, as its constraint ensures maximal accuracy. However, this same constraint becomes a disadvantage in batch editing scenarios, as it results in model collapse Gupta et al. (2024); Yang et al. (2024). The collapse occurs due to a sharp increase in model perplexity, likely caused by the compounding effects of sequential edits.

### v2

@@ 5.2 Neighbor-Assisted Model Editing Results
To evaluate neighbor-assisted model editing method, we only conducted experiments on COUNTERFACT due to data limitations of ZsRE explained in Appendix B. We perform iterative model editing with the modified neighbor loss, excluding collapsed settings (evaluation results in Table 3).
We observed consistently higher specificity across all iterations when using neighbor-assisted editing (denoted by NA_) compared to setups without it. This increase in specificity was accompanied by an overall improvement in score. Specifically, we observed gains of up to 6 percentage points in success accuracy and up to 34 percentage points in specificity accuracy (NA_MEMIT on GPT-2 XL). Although there was a slight decrease in generalization, the gain in specificity was more substantial, increasing the score. Additionally, because no prefix was added to the neighboring knowledge, we investigate its role in Appendix F.2. We found that adding prefixes led to a slightly higher overall score but reduced the specificity score.
Moreover, the proposed stopping criteria $|\Delta p_{k}|\leq\epsilon$, originally defined for iterative model editing, remain effective for neighbor-assisted model editing. We observed one exception in the case of LLaMA-2 (7B), where neighbor-assisted editing with PMET resulted in performance degradation. We attribute this to an increased tendency toward model collapse, as indicated by elevated perplexity (ME-PPL-50). Notably, LLaMA-2 (7B) was again the only model to exhibit such collapse behavior, reinforcing our earlier hypothesis that model-specific factors contribute to instability. However, identifying the precise training-related causes of this behavior requires deeper investigation, which we leave for future work.
@@
5.3 Analysis: How effective is the stopping criterion?
@@ 6 Related Work
In this work, we extensively discussed locate-and-edit model editing algorithms. KN (Dai et al., 2022) is a related method based on gradient-based neuron selection. In addition, there is a body of research that employs meta-learners to guide the parameter updates required for specific edits. For example, KE Cao et al. (2021) uses a hyper-network to update model parameters, while MEND Mitchell et al. (2022a) trains gradient-based, lightweight model editor networks. MALMEN Tan et al. (2024) builds upon MEND to address scalability challenges. Another line of research adds new knowledge without altering the model’s parameters. SERAC Mitchell et al. (2022b), GRACE Hartvigsen et al. (2023), and WISE Wang et al. (2024a) achieve this by employing additional memory to store new knowledge. A router network is then trained to decide whether to retrieve knowledge from the original model or the additional memory, ensuring the intended knowledge is accessed without modifying the model’s core parameters. We specifically focus on locate-and-edit model editing methods due to their effectiveness and efficiency in updating only the important parameters. Our proposed method introduces simple changes to existing techniques while still demonstrating effectiveness.
@@ 8 Limitations
Limited computational resources restricted us from experimenting with larger batch sizes and additional LLMs, such as GPT-NeoX (20B) and larger Llama-2 and Llama-3.1 models. We hypothesize that the experimental result trend will remain the same, and we leave the verification of this hypothesis for future work.
Appendix A Locate-and-Edit Algorithms
All locate-and-edit algorithms can be formulated under a unified two-stage framework consisting of an Optimization stage, where the target representation is computed, and a Spread stage, where the model weights are updated accordingly. As summarized in Table 1, this abstraction captures a wide range of existing model editing methods, and provides a foundation on which our proposed strategies—iterative and neighbor-assisted model editing—can be broadly applied. While some prior methods, such as AlphaEdit and EVOKE (LTI), specifically target the OverEdit problem, they do so using fundamentally different mechanisms from ours and do not generalize across algorithms. Notably, no existing work has directly addressed the UnderEdit challenge. To our knowledge, we are the first to propose a unified strategy that simultaneously mitigates both UnderEdit and OverEdit. Detailed descriptions of each algorithm are provided below.
As discussed in Section 2, model editing algorithms operate on the hypothesis that updating the final MLP parameters is sufficient to increase the likelihood of a new object $o^{*}$ over the original object $o$ when presented with a (subject, relation) pair $x=(s,r)$ as input to the LLM. Specifically, the final MLP weight matrix $W$ functions as a linear associative memory that stores a key-value mapping $[k,v]$ [^4] Distinct from key-value pairs in attention mechanismsMeng et al. (2023a). Here, the key encodes the last subject token, while the value represents the relation-object pair $(r,o)$ as a property of the subject $s$. Within the transformer architecture, the key corresponds to the output of the first fully connected MLP layer, whereas the value corresponds to the output of the second fully connected MLP layer $z$, as shown in Figure2. This key-value mapping $[k,v]$ is derived by computing the inner product between the key $k$ and the final MLP weight matrix $W$ as $Wk\approx v$.
@@ Appendix A Locate-and-Edit Algorithms
MEMIT
Meng et al. (2023b) operates under the assumption that factual edits can be made by modifying the final MLP weight matrix $W^{l}$ at each causal layer $l$ with a small update $\Delta^{l}$, yielding new weights $W^{l^{*}}=W^{l}+\Delta^{l}$. The goal is to insert $M\gg 1$ new key-value mappings $[K_{M},V_{M}]$ while preserving $E$ existing mappings $[K_{E},V_{E}]$, where $K_{E}=[k_{e}]_{e=1}^{E}$ and $V_{E}=[v_{e}]_{e=1}^{E}$ denote the pre-existing keys and values.
MEMIT formulates this as an optimization problem to find a transformation $\hat{W}$ that minimizes the sum of squared distances between transformed keys and their target values:
$\hat{W}$ $\triangleq\arg\min_{\hat{W}}\sum_{i=1}^{E+M}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}.$
This objective consists of two parts:
  • $\sum_{i=1}^{E}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}$: encourages the preservation of existing knowledge.
  • $\sum_{i=E+1}^{E+M}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}$: enforces the integration of new knowledge.
To compute the optimal update $\Delta^{l}$, MEMIT uses a closed-form solution derived from the residual matrix $R^{l}=V-WK$, where $K$ and $V$ stack the relevant key and value vectors across edits. Specifically,
$\Delta^{l}\leftarrow R^{l}K^{l\top}(C^{l}+K^{l}K^{l\top})^{-1},$
where $C^{l}$ is a regularization term proportional to the uncentered covariance of the pre-existing keys. This provides an analytical solution for $\Delta^{l}$, avoiding iterative optimization such as gradient descent. The resulting update is then applied as:
$W^{l^{*}}\leftarrow W^{l}+\Delta^{l}.$
In practice, the residual $R^{l}$ is computed using the hidden state $\bar{h}^{l_{c}}_{t}$ obtained in the Optimization stage. For full derivation and further implementation details, see Meng et al. (2023b).
@@
Li et al. (2024) shares the same optimization objective as MEMIT in the Spread stage but introduces a key difference in the Optimization stage. As shown in Figure 2, PMET searches for an ideal self-attention output $\bar{\textit{attn}}^{l_{c}}_{t}$ and an ideal MLP output $\bar{z}^{l_{c}}_{t}$. The core insight behind PMET is that the self-attention module captures generalizable patterns, while the MLP is more tightly coupled to fact-specific content. Therefore, PMET assumes that the contribution of self-attention to the hidden state $h^{l_{c}}_{t}$ is not necessary for editing factual knowledge.
Based on this assumption, PMET reconstructs a modified hidden state using only the ideal MLP output $\bar{z}^{l_{c}}_{t}$, effectively omitting the influence of attention. This ideal hidden state is then used to compute the residual matrix $R^{l}$, and the update $\Delta^{l}$ is computed using the same closed-form solution as MEMIT:
$\Delta^{l}\leftarrow R^{l}K^{l\top}(C^{l}+K^{l}K^{l\top})^{-1},$
followed by
$W^{l^{*}}\leftarrow W^{l}+\Delta^{l}.$
By decoupling attention from factual edits, PMET enables more precise updates that reduce unintended interference with unrelated knowledge. Empirically, this leads to improved edit success and generalization compared to MEMIT.
AlphaEdit
Fang et al. (2025) extends the locate-and-edit paradigm by explicitly constraining model updates to reduce interference with existing knowledge. While it follows a similar two-stage structure (Optimization and Spread), AlphaEdit introduces a novel use of null space projection (Wang et al., 2021) to isolate updates from directions associated with pre-existing knowledge.
During the Optimization stage, AlphaEdit computes the target hidden representation $\bar{h}^{l_{c}}_{t}$ in a manner similar to MEMIT. However, before computing the update $\Delta^{l}$, it projects the residual matrix $R^{l}=V-WK$ into the null space of the pre-existing keys $K_{E}$. This projection ensures that the update is orthogonal to directions associated with existing key-value mappings, thereby reducing the risk of OverEdit.
Formally, let $N^{l}$ be a projection matrix that spans the null space of $K_{E}$. AlphaEdit applies this projection to both the residual and key matrices:
$\Delta^{l}\leftarrow(N^{l}R^{l})(N^{l}K^{l})^{\top}\left(C^{l}+(N^{l}K^{l})(N^{l}K^{l})^{\top}\right)^{-1}.$
The resulting update is applied in the usual manner:
$W^{l^{*}}\leftarrow W^{l}+\Delta^{l}.$
By constraining updates to directions that are orthogonal to known information, AlphaEdit aims to improve specificity and mitigate OverEdit. However, as we show in Section 5, this constraint can limit editing flexibility, and iterative model editing can further improve AlphaEdit’s performance.
@@
Meng et al. (2023a) is the predecessor of MEMIT and adopts a more constrained approach to model editing. Unlike MEMIT, which apply updates across multiple causal layers, ROME assumes that new knowledge can be fully integrated into a single transformer layer $l_{*}$. It follows the same two-stage structure but focuses exclusively on editing one final MLP weight matrix $W^{l_{*}}$.
In the Optimization stage, ROME identifies an ideal MLP output $\bar{z}^{l_{*}}_{t}$ for a given input $(s,r)$. This output represents the desired value vector that should be produced by the edited layer for the edited subject. The Spread stage then computes an updated MLP weight matrix $\hat{W}^{l_{*}}$ that (1) preserves all existing knowledge and (2) exactly maps a new key $k_{*}$ to the target value $v_{*}$. Formally, the update is obtained by solving the following constrained optimization problem:
minimize $\left\|\hat{W}^{l_{*}}K_{E}-V_{E}\right\|^{2}$
subject to $\hat{W}^{l_{*}}k_{*}=v_{*},$
where $K_{E}$ and $V_{E}$ are matrices containing the keys and values for existing knowledge.
The resulting solution allows for a rank-preserving edit that satisfies the new constraint while minimizing distortion to previously stored mappings. For a complete derivation, see Meng et al. (2023a).
While ROME produces highly precise edits—especially for single-fact updates—it does not support true batch editing. [^5] Batch edits in ROME are executed sequentially, one fact at a time. This makes it prone to instability when applying many edits, as sequential updates can interfere destructively. Prior work (Gupta et al., 2024b; Yang et al., 2024) shows that ROME suffers from model collapse when used iteratively, due to sharp increases in perplexity arising from the accumulation of such updates.
R-ROME
Gupta et al. (2024a) builds on ROME with the explicit goal of resolving its tendency to collapse under sequential edits. While ROME applies a constrained rank-one update to a single MLP weight matrix $W^{l_{*}}$, it suffers from instability when edits are applied repeatedly. R-ROME attributes this collapse to how ROME computes the key vector $k$ used in the update rule.
In ROME, the key $k$ is directly computed from the current subject $s$, without considering broader contextual signals. R-ROME proposes a revised formulation in which the key $k_{*}$ is computed as an average over multiple neighboring prompts $x_{j}+s$, resulting in a more stable and robust key representation. This modification aligns the computation of the key $k_{*}$ and the value $v_{*}$, which are jointly used to derive the update.
The updated weight matrix is then computed via:
$\hat{W}=W+\Lambda_{*}(C^{-1}k_{*})^{\top},$
where $\Lambda_{*}$ scales the residual $v_{*}-Wk_{*}$ based on the adjusted key $k_{*}$. By unifying the key used in both the value computation and the update direction, R-ROME mitigates the destructive interference observed in ROME and reduces the likelihood of collapse from so-called “disabling edits”.
Empirically, R-ROME improves stability in sequential editing scenarios while retaining the precision benefits of ROME. However, like ROME, it remains limited to editing one fact at a time and does not natively support batch updates.
EMMET
Gupta et al. (2024c) unifies the ROME and MEMIT families of model editing by showing that both optimize a common preservation-memorization objective. While ROME performs single edits using strict equality constraints and MEMIT enables batched edits via a least-squares formulation, EMMET generalizes both by introducing equality-constrained batch editing. This allows EMMET to combine the precision of ROME with the scalability of MEMIT.
Like other methods in the locate-and-edit family, EMMET follows the two-stage editing framework. In the Optimization stage, it identifies the target value vectors $V_{E}$ corresponding to the new facts to be inserted. In the Spread stage, it solves a constrained optimization problem to update the MLP weight matrix $W_{0}$ to a new matrix $\hat{W}$ that satisfies two goals: (1) preserving the projections of a set of pre-existing keys $K_{0}$, and (2) exactly mapping a batch of new keys $K_{E}$ to their corresponding values $V_{E}$. Formally:
[…]
$\quad\hat{W}k^{(e)}_{i}=v^{(e)}_{i}\quad\forall i\in[1,E],$
where the first term enforces knowledge preservation and the constraints enforce exact memorization of the new facts.
This preservation-memorization objective is solved using Lagrange multipliers, yielding the closed-form update:
$\hat{W}=W_{0}+(V_{E}-W_{0}K_{E})(K_{E}^{\top}C_{0}^{-1}K_{E})^{-1}K_{E}^{\top}C_{0}^{-1},$
where $C_{0}=K_{0}K_{0}^{\top}$ is the uncentered covariance matrix of the preserved keys.
By design, EMMET enables high-precision edits with support for batch sizes up to 10,000, effectively bridging the capabilities of ROME and MEMIT within a unified formulation. Empirically, it achieves performance comparable to MEMIT while retaining the theoretical rigor of equality-constrained updates. As such, EMMET offers a principled and scalable approach to batch model editing that fits cleanly into the Optimization/Spread framework.
ENCORE
Gupta et al. (2025) addresses the long-term instability of locate-and-edit methods under large-scale sequential editing. Prior work has shown that repeated edits with methods like MEMIT and ROME lead to overfitting on edited facts and model collapse due to uncontrolled growth in the norm of updated weights. ENCORE introduces two key enhancements—Most-Probable Early Stopping (MPES) and a norm constraint—to enable robust, long-horizon editing.
In the Optimization stage, ENCORE adopts MPES, an early stopping strategy that halts optimization once the target object becomes the most probable prediction across all optimization queries. This prevents overfitting by stopping the editing process before excessive memorization occurs, similar to how early stopping in training halts based on validation loss.
In the Spread stage, ENCORE augments the standard preservation-memorization objective with a Frobenius norm penalty that discourages large deviations from the original weights:
$\mathcal{L}(\hat{W})=$ $\;\;\lambda_{p}\|\hat{W}K_{0}-W_{0}K_{0}\|^{2}$
$+\|\hat{W}K_{1}-V_{1}\|^{2}$
$+\lambda_{n}\|\hat{W}-W_{0}\|_{F}^{2},$
where the first two terms represent knowledge preservation and memorization (as in MEMIT), and the third term explicitly controls norm growth. This yields a closed-form solution:
$\hat{W}=W_{0}+(V_{1}-W_{0}K_{1})$ $K_{1}^{\top}\Big{(}\lambda_{p}K_{0}K_{0}^{\top}$
$+K_{1}K_{1}^{\top}+\lambda_{n}I\Big{)}^{-1}.$
Like other methods in this space, ENCORE cleanly fits into the two-stage locate-and-edit framework, using MPES for target identification in Optimization and a norm-constrained update formulation in Spread.
EVOKE (LTI)
Zhang et al. (2025) introduces a plug-and-play optimization strategy called Learn the Inference (LTI) to mitigate OverEdit in complex reasoning tasks such as multi-hop inference. While prior locate-and-edit methods like ROME and MEMIT often overfit to edit targets, assigning disproportionately high probabilities to them, EVOKE attributes this Editing Overfit to the strong coupling between the edit prompt and the target object. To address this, LTI regularizes the optimization process by incorporating auxiliary constraints inspired by how unedited LLMs recall knowledge via in-context learning.
EVOKE operates within the standard two-stage editing framework. In the Optimization stage, it modifies the optimization objective used to compute the target value vector $v^{*}$, introducing three additional constraints:
  • Subject Representation Constraint (SRC): aligns the intermediate representation of the subject token between the edited and unedited models to avoid overfitting on narrow context.
  • Output Distribution Constraint (ODC): matches the output distributions of the edited and unedited models to preserve global behavior.
  • New Knowledge Constraint (NKC): ensures the edited model correctly predicts the new target object across randomly prefixed contexts.
These constraints are jointly optimized via a weighted objective:
$\mathcal{L}=\lambda\,\mathcal{L}_{\mathrm{SRC}}+\beta\,\mathcal{L}_{\mathrm{ODC}}+\alpha\,\mathcal{L}_{\mathrm{NKC}},$
where $\lambda,\beta,\alpha$ are tunable weights. Once the optimal residual vector $h$ is learned, the final update in the Spread stage follows the standard weight update (e.g., as used in ROME).
While EVOKE aims to mitigate OverEdit through more informed optimization, its approach differs from our neighbor-assisted editing strategy in key ways. EVOKE constrains the model’s internal representations by comparing to unedited inference with prepended context, whereas we use neighboring samples that share the same relation to guide editing. Additionally, EVOKE perturbs the value vector $v^{*}$ while holding the subject fixed and varying the relation; in contrast, we vary the subject and fix the relation to ensure the original object remains unchanged. EVOKE also uses significantly more prompts per edit sample, which may affect scalability.
Despite these differences, EVOKE still adheres to the two-stage locate-and-edit framework. As such, our proposed iterative and neighbor-assisted methods are compatible with it and could further enhance its performance by addressing UnderEdit and reinforcing relation-consistent specificity.